# A Laboratory Experiment on Using Different Financial-Incentivization Schemes in Software-Engineering Experimentation

### v5

@@ 3.2 Laboratory Experiment
Goal. After eliciting which payoff functions are used and preferred in practice, we will conducted our actual experiment to measure the impact of different payoff functions in software-engineering experiments. We focused on code reviews and bug identification in this experiment, since these are typical tasks in software engineering that also involve different types of incentives. So, we aimed to support software-engineering researchers by identifying which payoff functions can be used to improve the validity of experiments.
Treatments. As motivated, we aimed to compare four treatments to reflect different payoff functions that stemmed from our survey and established research. While we are were able to define the payoff functions for the “No Performance Incentives Treatment” (NPIT) and “Open-Source Incentives Treatment” (OSIT) in advance, we needed data from our survey to proceed with the “MP Incentives Treatment” (MPIT) and “MA Incentives Treatment” (MAIT). However, we can did a priori describe the method we would use to derive the payoff functions for those treatments. Note that some treatments may could yield the same payoff function (i.e., NPIT, MAIT, and MPIT). It is unlikely that all three payoff functions will would be identical, but we will merged those that are were (i.e., NPIT and MAIT) and reduced the number of treatments accordingly (see Table 2 for the variable names):
@@
For NPIT, we provided a fixed payment (i.e., 10 €) that will be was payed out at the end of an experimental session. So, this treatment mimics a participation fee for experiments or fixed wages for the real world. Consequently, the payoff is independent of a participant’s actual performance. Overall, the payoff function ($PF$) for this treatment is:
@@
Again, this treatment does not depend on our survey results, but builds on findings from the software-engineering literature on the motivation of open-source developers (Gerosa et al., 2021; Hertel et al., 2003; Hars and Ou, 2002; Ye and Kishida, 2003; Huang et al., 2021). We remark that we focused particularly on those developers that do not receive payments (e.g., as wages or bug bounties), but work for free. In a simplified, economics perspective, such developers still act within a conceptual cost-benefit framework (i.e., they perceive to obtain a benefit from working on the software). We build built on the induced-value method (Weimann and Brosig-Koch, 2019) from experimental economics to mimic this cost-benefit framework with financial incentives to derive the OSIT treatment. Besides a participation fee, we will involved a performance-based reward for correctly identifying all bugs to resemble goal-oriented incentives (e.g., personal fulfillment of achieving a goal or supporting open-source projects). Furthermore, we considered the opportunity costs of working on open-source software (i.e., less time for other projects and additional effort for performing a number of checks). Overall, the payoff function ($PF$) for this treatment is:
$PF_{\text{\textst{NPIT}}OSIT}=payment_{fix}+reward_{complete}-time\cdot penalty_{time}-checks\cdot penalty_{checks}$
@@
Using our survey results, we will be able the could identify a payoff function that represents what is mostly applied in practice. We will would then derive a payoff function as explained in Section 3.1. However, we found that the survey results led to the same function as for NPIT, which is why we did not use a distinct MAIT in our actual experiment.
@@
We will used the same method we used for MAIT to define a payoff function for MPIT. In this case, the developers preferred a fixed payment with an additional quality reward that is based on their organization’s performance:
@@
We used the same code-review example for all treatments to keep the complexity of the problem constant. For all treatments, we will calibrated the payoff function so that the expected payoff for each participant in and between treatments is was approximately the same (i.e., around 10 €). Implementing similar expected payoffs avoids unfairness between treatments, and ensures that performance differences are caused by different incentive schemes and not the total size of the payoff. After the treatment, we will gathered demographic data from the participants (e.g., age, gender) and asked for any concerns or feedback. We estimated that each sessions of the experiment would take 45 minutes.
@@
Code Example. We selected and adapted three different Java code examples (i.e., limited calculator, sorting and searching, a Stack), the first from the learning platform LeetCode [^4] https://leetcode.com and the other two from the “The Algorithms” GitHub repository. [^5] https://github.com/TheAlgorithms/Java To create buggy examples, we injected three bugs into each code example by using mutation operators (Jia and Harman, 2011). Note that we partly reworked the examples to make them more suitable for our experiment (e.g., combining searching and sorting), added comments at the top of each example explaining its general purpose, and kept other comments (potentially adapted) as well as identifier names to improve the realism. We aimed to limit the time of the experiment to avoid fatigue and actually allow for a laboratory setting, and thus decided to use only one example. To select the most suitable subject system for our experiment, we performed a pilot study in which we measured the time and performance of the participants. In detail, we asked one M.Sc. student from the University of Glasgow who has worked as a software practitioner in industry and four PhD students from the University of Zurich to perform the code reviews on the buggy examples. Overall, each example was reviewed by three of these participants. Our results indicated that the sorting and searching example would be most feasible (i.e., $\approx$12 min., 4/9 bugs correctly identified, 5 false positives), considering that the task should neither be too easy nor to hard, the background of the pilot’s participants and the potential participants for our experiment, as well as the examples’ quality. The other two examples seemed too large or complicated (i.e., $\approx$14 min., 2/9 bugs; 4 false positives; $\approx$8 min., 5/9 bugs, 8 false positives), which is why we decided to use the sorting and searching example (available in our artifacts).2 We remark that none of the participants from this pilot study will be was involved in our actual experiment. In Figure 1, we display a screenshot of the sorting and searching code example as we showed will show it to the participants in the lab.
@@
Figure 1: Screenshot of the code example as it will be shown we showed it to the participants. The checkboxes in front of each line allowed the participants to check buggy lines of code. Note that we will did not show the comments indicating the implemented bugs (i.e., in lines 16, 21, and 38). The blue boxes (not displayed for participants) indicate the Areas of Interest (AOIs) that we used for the eye-tracking analysis.
@@
Besides analyzing these hypotheses, we will also compared the behavior (e.g. risk taking) and performance between all groups to understand which incentives have what impact. Moreover, we will used eye trackers to explore fixation counts, fixation lengths, and return fixations. This will allow allowed us to obtain a deeper understanding of the search and evaluation processes during code reviews. Also, it enablesd us to investigate potential differences in eye movements depending on the incentivization. More precisely, we intended to follow similar studies from software engineering Abid et al. (2019) to explore how our participants read the source code, for instance, do did they focus on the actually buggy code, what lines are were they reading more often, or which code elements do did they focus on to explore bugs? We will report all findings from the eye-tracking data as exploratory outcomes. The eye-tracking data is preprocessed by the firmware of Tobii (Version 1.181.37603) using the Tobii I-VT (fixation) filter.
@@
Metrics. The performance of our participants is was primarily depending on their correctness in identifying bugs during the code review. Since this can be expressed as confusion matrices, we decided to implement the F1-score (defined as $\frac{2TP}{2TP+FP+FN}$) as the only outcome measure to evaluate our hypotheses. For our experiment, true positives (TP) refer to the correctly identified bugs, false positives (FP) refer to the locations marked as buggy that are actually correct, and false negatives (FN) refer to the undetected bugs. Note that our participants will were not be aware of this metric (they will only know knew about the payoff function) to avoid biases, and any decision based on the payoff function will be are reflected by the F1-score (e.g., taking more risks due to missing penalties under NPIT). So, this metric allowsed us to compare the performances of our participants between treatments considering that they motivate different behaviors, which allowsed us to test our hypotheses. In summary, our dependent variable is was the F1-score, our independent variable is was the payoff function, and we will collected additional data via a post experimental survey (e.g., experience, gender, age, stress) as well as eye-tracking data for exploratory analyses.
@@
Experimental Design. Fundamentally, we will employ a 4x1 design (alternatively, if the survey indicates no differences between MPIT and MAIT, a 3x1 design). Fundamentally, we planned to employ a 4x1 design. However, since we merged the treatments NPIT and MAIT after our survey, we ended up with a 3x1 design). For each treatment, we only changed the payoff function. We will allocated participants to their treatment at random, without anyone repeating the experiment in another treatment. On-site, we can could execute the experiment at the experimental laboratory MaXLab of the Otto-von-Guericke University using a standardized experimental environment. We will employed a between-subject design measuring the participants’ performance and measured the eye movement of four participants (restricted by number of devices) in each session using eye trackers (60 Hz Tobii Pro Nano H). Note that we will could identify any impact wearing eye-trackers may have had on our participants during our analysis. However, it is not likely that they will have had an impact, because this type of eye trackers is mounted to the screen and barely noticeable, not a helmet the participants have to wear. The procedure for each session is was as follows:
@@
After the participants of a session entered the laboratory, they are were randomly allocated to working stations with the experimental environment installed. Moreover, four of them are were randomly selected for using eye trackers. To this end, we will already stated in the invitation that eye tracking is would be involved in the experiment. If a participant nonetheless disagreesd to participate using eye trackers, we will excluded them from the experiment to avoid selection bias. Once all participants are were at their places, the experimenter begins began the experiment. The participants received general information about the experiment (e.g., welcoming text), information about the task at hand (code review), an explanation on how to enter data (e.g., check box), and the definition of their payoff function for the experiment (with some examples).
@@
All participants received the code example with the task to identify any bugs within it. Note that the participants will were not be aware of the precise number of bugs in the code. Instead, a message will explained that the code does not behave as expected when it is executed. At the end of the task, we can could have incorporated unpaid overtime as a payment component by asking participants to stay for five more minutes to work on the task.
@@
After the experiment, we will asked our participants a number of demographic questions (i.e., gender, age, study program, study term, programming experience). We will further apply applied the distress subscale of the Short Stress State Questionnaire (Helton, 2004) to measure arousal and stress of the participants. Eliciting such data on demographics and arousal will enabled us to identify potential confounding parameters.
@@
After we have collected all the data, we provided information about their performance and payoff to the participants by displaying them on their screen. We will payed out these earnings privately in a separate room in cash immediately afterwards.
Analysis Plan. To analyze our data, we will employed the following steps:
@@
The experimental environment storesd raw data in CSV files. We do did not plan to remove any outliers or data unless we would identify a specific reason for which we would believe the data would could be invalid, which involvesd primarily two cases. First, it may happen have happened that the eye-movement recordings of a participant have a low quality (i.e., $<$70 % gaze sample). Gaze sample is defined as the percentage of the time that the eyes are correctly detected. Since we used eye tracking only for exploratory analyses, we we will not replace subjects would not have replaced participants just because the calibration was not good enough. Moreover, the participants will not be were not aware of the quality and could simply continue with the actual experiment. However, we will excluded their eye-tracking data from our exploratory analysis. Second, we will would have excluded participants if they violated the terms of conduct of the laboratory. While this case is was unlikely, we will try would have tried to replace these participants to achieve the desired sample sizes Yet, we would do this before data cleaning. Fortunately, neither of such cases occured.
@@
We will present used descriptive statistics for the demographic, dependent, and independent variables for each treatment by, reporting means and standard deviations of the respective variables.
@@
Since sole statistical testing is often subject to misinterpretation and not recommended (Wasserstein and Lazar, 2016; Wasserstein et al., 2019; Amrhein et al., 2019), we will focused on describing our observations. For this purpose, we willstarted with reporting the results we obtained, plotting suitable visualizations, and identifying patterns within these. The statistical tests will helped us to improve our confidence in these observations.
@@
For our analysis, we In total, this leads to 6 tests, or to 3 tests if our survey indicates that two treatments are mostly identical. focused on performance (i.e., F1 score). We will first checked whether the assumptions required for parametric tests (e.g., normality) are fulfilled, and if not proceeded with non-parametric tests (i.e., Wilcoxon-Mann-Whitney test). Since we are were interested in all possible differences between the four three treatments, we have had to conduct all pairwise treatment tests. In total, this leads to 6 tests, or to 3 tests if our survey indicates that two treatments are mostly identical. In total, this led to three tests. For the significance analyses, we will apply applied a significance level of $p<0.05$ and corrected for multiple hypotheses testing using the Holm-Bonferroni method. Though the share of participants who will used eye trackers will be was constant among all treatments, and thus should not affect treatment effects, we further checked whether the presence of eye trackers affected performance. To increase the statistical robustness, we will also conducted a regression analysis using the treatments as categorical variables and NPIT as base. As exogenous variables, we included: age, gender, experience, and arousal of the participants. In contrast to the preregistered tests, we will discuss these results as exploratory outcomes.
Based on these steps, we will obtained a detailed understanding of how different incetivization schemes impact the performance of software developers during code review.
@@ 4.2 Experiment
According to our registered report, we focused on the F1 score as the measure of participants’ performance. As our experimental data does not fulfill the assumptions for a parametric test (Shapiro-Wilk test, NPIT: p-value $<$ 0.01, OSIT: p-value $<$ 0.01, MPIT: p-value $<$ 0.01), we proceeded with the Wilcoxon-Mann-Whitney test for our statistical tests. To investigate H1 (cf. Table 1), we compared NPIT with OSIT and MPIT, respectively. Despite the notable differences in the F1 scores (0.26 vs 0.16 and 0.15), our statistical tests indicate no significant result (NPIT-OSIT: p-value = 0.896, padjusted = 1, NPIT-MPIT: p-value = 0.923, padjusted = 1), which is in large part due to our hypothesis stating that participants would perform better when performance incentives are in place. Instead, we see indications for the opposite. This is a surprising result, and we will provide some insights on possible explanations in the exploratory analysis. With respect to the two performance-dependent treatments (MPIT, OSIT), we also see no significant differences with respect to the F1 score (p-value = 0.796, padjusted = 1).
As the last step of our preregistered analysis plan, we conducted a regression analysis. The results of the Tobit regression with limits at 0 and 1 (cf. Table 7) mostly confirm our previous findings (performance in NPIT is insignificantly better than in OSIT and MPIT). Yet, adding a parameter (completion Time) that we did not preregister in model (3) indicates the importance of the completion time on the F1 scores. The longer the participants stayed in the experiment, the higher was their F1 score. We will address the topic of completion time in more detail in the following exploratory analysis.
@@
Dependent variable: F1 exploratory (1) (2) (3) treatmentOSIT --0.171 (0.132) --0.144 (0.138) --0.054 (0.136) treatmentMPIT --0.134 (0.128) --0.146 (0.137) --0.208 (0.134) age --0.004 (0.014) --0.010 (0.013) genderWoman 0.176 (0.122) 0.175 (0.116) programmingExperience --0.003 (0.034) --0.016 (0.033) engagement 0.018 (0.043) 0.042 (0.042) distress --0.042 (0.048) --0.060 (0.046) worry 0.005 (0.042) --0.005 (0.041) completionTime ${}^{**}$ (0.006) logSigma ${}^{***}$ (0.141) ${}^{***}$ (0.141) ${}^{***}$ (0.140) constant 0.139 (0.094) 0.213 (0.417) 0.144 (0.399) $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01
@@
Eye-Tracking Analysis. Approximately half of our participants in every treatment conducted the experiment with eye trackers. We can see no evidence that eye-tracking changed their performance (Wilcoxon-Mann-Whitney test, NPIT: p-value = 0.702 padjusted = 1, OSIT: p-value = 0.277, padjusted = 0.831, MPIT: p-value = 0.535, padjusted = 1). After evaluating the quality of the eye-tracking data, we had to exclude seven of 31 observations due to (1) low gaze detection ($<$70 %) during the whole timespan or (2) high validation accuracy ($>$1.5°) and high validation precision ($>$1°) during the eye tracking calibration. This left us with 7/7/10 observations in NPIT/OSIT/MPIT, respectively. Still, the eye-tracking data provides us with valuable information on the participants’ behavior.

### v7

@@ 3.2 Laboratory Experiment
Goal. After eliciting which payoff functions are used and preferred in practice, we conducted our actual experiment to measure the impact of different payoff functions in software-engineering experiments. We focused on code reviews and bug identification in this experiment, since these are typical tasks in software engineering that also involve different types of incentives. So, we aimed to support software-engineering researchers by identifying which payoff functions can be used to improve the validity of experiments.
Treatments. As motivated, we aimed to compare four treatments to reflect different payoff functions that stemmed from our survey and established research. While we were able to define the payoff functions for the “No Performance Incentives Treatment” (NPIT) and “Open-Source Incentives Treatment” (OSIT) in advance, we needed data from our survey to proceed with the “MP Incentives Treatment” (MPIT) and “MA Incentives Treatment” (MAIT). However, we did a priori describe the method we would use to derive the payoff functions for those treatments. Note that some treatments could yield the same payoff function (i.e., NPIT, MAIT, and MPIT). It is unlikely that all three payoff functions would be identical, but we merged those that were (i.e., NPIT and MAIT) and reduced the number of treatments accordingly (see Table 2 for the variable names):
@@
For NPIT, we provided a fixed payment (i.e., 10 €) that was payed out at the end of an experimental session. So, this treatment mimics a participation fee for experiments or fixed wages for the real world. Consequently, the payoff is independent of a participant’s actual performance. Overall, the payoff function ($PF$) for this treatment is:
@@
Again, this treatment does not depend on our survey results, but builds on findings from the software-engineering literature on the motivation of open-source developers (Gerosa et al., 2021; Hertel et al., 2003; Hars and Ou, 2002; Ye and Kishida, 2003; Huang et al., 2021). We remark that we focused particularly on those developers that do not receive payments (e.g., as wages or bug bounties), but work for free. In a simplified, economics perspective, such developers still act within a conceptual cost-benefit framework (i.e., they perceive to obtain a benefit from working on the software). We built on the induced-value method (Weimann and Brosig-Koch, 2019) from experimental economics to mimic this cost-benefit framework with financial incentives to derive the OSIT treatment. Besides a participation fee, we involved a performance-based reward for correctly identifying all bugs to resemble goal-oriented incentives (e.g., personal fulfillment of achieving a goal or supporting open-source projects). Furthermore, we considered the opportunity costs of working on open-source software (i.e., less time for other projects and additional effort for performing a number of checks). Overall, the payoff function ($PF$) for this treatment is:
$PF_{\text{}OSIT}=payment_{fix}+reward_{complete}-time\cdot penalty_{time}-checks\cdot penalty_{checks}$
@@
Using our survey results, we could identify a payoff function that represents what is mostly applied in practice. We would then derive a payoff function as explained in Section 3.1. However, we found that the survey results led to the same function as for NPIT, which is why we did not use a distinct MAIT in our actual experiment.
@@
We used the same method we used for MAIT to define a payoff function for MPIT. In this case, the developers preferred a fixed payment with an additional quality reward that is based on their organization’s performance:
@@
We used the same code-review example for all treatments to keep the complexity of the problem constant. For all treatments, we calibrated the payoff function so that the expected payoff for each participant in and between treatments was approximately the same (i.e., around 10 €). Implementing similar expected payoffs avoids unfairness between treatments, and ensures that performance differences are caused by different incentive schemes and not the total size of the payoff. After the treatment, we gathered demographic data from the participants (e.g., age, gender) and asked for any concerns or feedback. We estimated that each session of the experiment would take 45 minutes.
@@
Code Example. We selected and adapted three different Java code examples (i.e., limited calculator, sorting and searching, a Stack), the first from the learning platform LeetCode [^4] https://leetcode.com and the other two from the “The Algorithms” GitHub repository. [^5] https://github.com/TheAlgorithms/Java To create buggy examples, we injected three bugs into each code example by using mutation operators (Jia and Harman, 2011). Note that we partly reworked the examples to make them more suitable for our experiment (e.g., combining searching and sorting), added comments at the top of each example explaining its general purpose, and kept other comments (potentially adapted) as well as identifier names to improve the realism. We aimed to limit the time of the experiment to avoid fatigue and actually allow for a laboratory setting, and thus decided to use only one example. To select the most suitable subject system for our experiment, we performed a pilot study in which we measured the time and performance of the participants. In detail, we asked one M.Sc. student from the University of Glasgow who has worked as a software practitioner in industry and four PhD students from the University of Zurich to perform the code reviews on the buggy examples. Overall, each example was reviewed by three of these participants. Our results indicated that the sorting and searching example would be most feasible (i.e., $\approx$12 min., 4/9 bugs correctly identified, 5 false positives), considering that the task should neither be too easy nor to hard, the background of the pilot’s participants and the potential participants for our experiment, as well as the examples’ quality. The other two examples seemed too large or complicated (i.e., $\approx$14 min., 2/9 bugs; 4 false positives; $\approx$8 min., 5/9 bugs, 8 false positives), which is why we decided to use the sorting and searching example (available in our artifacts).2 We remark that none of the participants from this pilot study was involved in our actual experiment. In Figure 1, we display a screenshot of the sorting and searching code example we showed to the participants in the lab.
@@
Figure 1: Screenshot of the code example as we showed it to the participants. The checkboxes in front of each line allowed the participants to check buggy lines of code. Note that we did not show the comments indicating the implemented bugs (i.e., in lines 16, 21, and 38). The blue boxes (not displayed to participants) indicate the Areas of Interest (AOIs) that we used for the eye-tracking analysis.
@@
Besides analyzing these hypotheses, we also compared the behavior (e.g. risk taking) and performance between all groups to understand which incentives have what impact. Moreover, we used eye trackers to explore fixation counts, fixation lengths, and return fixations. This allowed us to obtain a deeper understanding of the search and evaluation processes during code reviews. Also, it enabled us to investigate potential differences in eye movements depending on the incentivization. More precisely, we intended to follow similar studies from software engineering Abid et al. (2019) to explore how our participants read the source code, for instance, did they focus on the actually buggy code, what lines were they reading more often, or which code elements did they focus on to explore bugs? We report all findings from the eye-tracking data as exploratory outcomes. The eye-tracking data is preprocessed by the firmware of Tobii (Version 1.181.37603) using the Tobii I-VT (fixation) filter.
@@
Metrics. The performance of our participants was primarily depending on their correctness in identifying bugs during the code review. Since this can be expressed as confusion matrices, we decided to implement the F1-score (defined as $\frac{2TP}{2TP+FP+FN}$) as the only outcome measure to evaluate our hypotheses. For our experiment, true positives (TP) refer to the correctly identified bugs, false positives (FP) refer to the locations marked as buggy that are actually correct, and false negatives (FN) refer to the undetected bugs. Note that our participants were not aware of this metric (they only knew about the payoff function) to avoid biases, and any decision based on the payoff function are reflected by the F1-score (e.g., taking more risks due to missing penalties under NPIT). So, this metric allowed us to compare the performances of our participants between treatments considering that they motivate different behaviors, which allowed us to test our hypotheses. In summary, our dependent variable was the F1-score, our independent variable was the payoff function, and we collected additional data via a post experimental survey (e.g., experience, gender, age, stress) as well as eye-tracking data for exploratory analyses.
@@
Experimental Design. Fundamentally, we planned to employ a 4x1 design. However, since we merged the treatments NPIT and MAIT after our survey, we ended up with a 3x1 design). For each treatment, we only changed the payoff function. We allocated participants to their treatment at random, without anyone repeating the experiment in another treatment. On-site, we could execute the experiment at the experimental laboratory MaXLab of the Otto-von-Guericke University using a standardized experimental environment. We employed a between-subject design measuring the participants’ performance and measured the eye movement of four participants (restricted by number of devices) in each session using eye trackers (60 Hz Tobii Pro Nano H). Note that we could identify any impact wearing eye-trackers may have had on our participants during our analysis. However, it is not likely that they had an impact, because this type of eye trackers is mounted to the screen and barely noticeable, not a helmet the participants have to wear. The procedure for each session was as follows:
@@
After the participants of a session entered the laboratory, they were randomly allocated to working stations with the experimental environment installed. Moreover, four of them were randomly selected for using eye trackers. To this end, we already stated in the invitation that eye tracking would be involved in the experiment. If a participant nonetheless disagreed to participate using eye trackers, we excluded them from the experiment to avoid selection bias. Once all participants were at their places, the experimenter began the experiment. The participants received general information about the experiment (e.g., welcoming text), information about the task at hand (code review), an explanation on how to enter data (e.g., check box), and the definition of their payoff function for the experiment (with some examples).
@@
All participants received the code example with the task to identify any bugs within it. Note that the participants were not aware of the precise number of bugs in the code. Instead, a message explained that the code does not behave as expected when it is executed. At the end of the task, we could have incorporated unpaid overtime as a payment component by asking participants to stay for five more minutes to work on the task.
@@
After the experiment, we asked our participants a number of demographic questions (i.e., gender, age, study program, study term, programming experience). We further applied the distress subscale of the Short Stress State Questionnaire (Helton, 2004) to measure arousal and stress of the participants. Eliciting such data on demographics and arousal enabled us to identify potential confounding parameters.
@@
After we collected all the data, we provided information about their performance and payoff to the participants by displaying them on their screen. We payed out these earnings privately in a separate room in cash immediately afterwards.
Analysis Plan. To analyze our data, we employed the following steps:
@@
The experimental environment stored raw data in CSV files. We did not plan to remove any outliers or data unless we would identify a specific reason for which we would believe the data could be invalid, which involved primarily two cases. First, it may have happened that the eye-movement recordings of a participant have a low quality (i.e., $<$70 % gaze sample). Gaze sample is defined as the percentage of the time that the eyes are correctly detected. Since we used eye tracking only for exploratory analyses, we would not have replaced participants just because the calibration was not good enough. Moreover, the participants were not aware of the quality and could simply continue with the actual experiment. However, we excluded their eye-tracking data from our exploratory analysis. Second, we would have excluded participants if they violated the terms of conduct of the laboratory. While this case was unlikely, we would have tried to replace these participants to achieve the desired sample sizes before data cleaning. Fortunately, neither of such cases occured.
@@
We used descriptive statistics for the demographic, dependent, and independent variables for each treatment , reporting means and standard deviations of the respective variables.
@@
Since sole statistical testing is often subject to misinterpretation and not recommended (Wasserstein and Lazar, 2016; Wasserstein et al., 2019; Amrhein et al., 2019), we focused on describing our observations. For this purpose, we started with reporting the results we obtained, plotting suitable visualizations, and identifying patterns within these. The statistical tests helped us to improve our confidence in these observations.
@@
For our analysis, we focused on performance (i.e., F1 score). We first checked whether the assumptions required for parametric tests (e.g., normality) are fulfilled, and if not proceeded with non-parametric tests (i.e., Wilcoxon-Mann-Whitney test). Since we were interested in all possible differences between the three treatments, we had to conduct all pairwise treatment tests. For the significance analyses, we applied a significance level of $p<0.05$ and corrected for multiple hypotheses testing using the Holm-Bonferroni method. Although the share of participants who used eye trackers was constant among all treatments, and thus should not affect treatment effects, we further checked whether the presence of eye trackers affected performance. To increase the statistical robustness, we also conducted a regression analysis using the treatments as categorical variables and NPIT as base. As exogenous variables, we included: age, gender, experience, and arousal of the participants. In contrast to the preregistered tests, we discuss these results as exploratory outcomes.
Based on these steps, we obtained a detailed understanding of how different incetivization schemes impact the performance of software developers during code review.
@@ 4.2 Experiment
According to our registered report, we focused on the F1 score as the measure of participants’ performance. As our experimental data does not fulfill the assumptions for a parametric test (Shapiro-Wilk test, NPIT: p-value $<$ 0.01, OSIT: p-value $<$ 0.01, MPIT: p-value $<$ 0.01), we proceeded with the Wilcoxon-Mann-Whitney test for our statistical tests. Adjusted p-values (padjusted) stem from the Holm-Bonferroni correction. To investigate H1 (cf. Table 1), we compared NPIT with OSIT and MPIT, respectively. Despite the notable differences in the F1 scores (0.26 vs 0.16 and 0.15), our statistical tests indicate no significant result (NPIT-OSIT: p-value = 0.896, padjusted $>$0.99)= 1, NPIT-MPIT: p-value = 0.923, padjusted $>$0.99)= 1), which is in large part due to our hypothesis stating that participants would perform better when performance incentives are in place. Instead, we see indications for the opposite. This is a surprising result, and we will provide some insights on possible explanations in the exploratory analysis. With respect to the two performance-dependent treatments (MPIT, OSIT), we also see no significant differences with respect to the F1 score (p-value = 0.796, padjusted $>$0.99)= 1).
As the last step of our preregistered analysis plan, we conducted a regression analysis. The results of the Tobit regression with limits at 0 and 1 (cf. Table 7) mostly confirm our previous findings (performance in NPIT is innon-significantly better than in OSIT and MPIT). Yet, adding a parameter (completion Time) that we did not preregister in model (3) indicates the importance of the completion time on the F1 scores. The longer the participants stayed in the experiment, the higher was their F1 score. We will address the topic of completion time in more detail in the following exploratory analysis.
@@
Dependent variable: F1 exploratory (1) (2) (3) treatmentOSIT --0.171 (0.132) --0.144 (0.138) --0.054 (0.136) treatmentMPIT --0.134 (0.128) --0.146 (0.137) --0.208 (0.134) age --0.004 (0.014) --0.010 (0.013) genderWoman 0.176 (0.122) 0.175 (0.116) programmingExperience --0.003 (0.034) --0.016 (0.033) engagement 0.018 (0.043) 0.042 (0.042) distress --0.042 (0.048) --0.060 (0.046) worry 0.005 (0.042) --0.005 (0.041) completionTime 0.016∗∗ (0.006) logSigma --0.927∗∗∗ (0.141) --0.955∗∗∗ (0.141) --1.012∗∗∗ (0.140) constant 0.139 (0.094) 0.213 (0.417) 0.144 (0.399) p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
@@
Eye-Tracking Analysis. Approximately half of our participants in every treatment conducted the experiment with eye trackers. We can see no evidence that eye-tracking changed their performance (Wilcoxon-Mann-Whitney test, NPIT: p-value = 0.702 padjusted $>$0.99)= 1), OSIT: p-value = 0.277, padjusted = 0.831, MPIT: p-value = 0.535, padjusted $>$0.99)= 1). After evaluating the quality of the eye-tracking data, we had to exclude seven of 31 observations due to (1) low gaze detection ($<$70 %) during the whole timespan or (2) high validation accuracy ($>$1.5°) and high validation precision ($>$1°) during the eye tracking calibration. This left us with 7/7/10 observations in NPIT/OSIT/MPIT, respectively. Still, the eye-tracking data provides us with valuable information on the participants’ behavior.